OpenStudy (anonymous):

how to find an irrational number? is -1.2(countinues) an irrational number?

OpenStudy (mertsj):

1.2222222........is not irrational because it repeats.

OpenStudy (anonymous):

so it is rational?

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

1.222222222... = 11/9 so it is rational

OpenStudy (anonymous):

irrationals cannot be written as p/q where p and q are integers

OpenStudy (anonymous):

thank you everyone :))

OpenStudy (anonymous):

If a terminating or non-terminating decimal can be represented as p/q form where p and q are integers and \(q \neq 0 \) Then that number is a rational number.

OpenStudy (anonymous):

to generate some irrational numbers , let k be any rational, then k + sqrt(2) is irrational Proof: let k = p/q assume k + sqrt(2) = m/n (assume it is rational) p/q + sqrt(2) = m/n sqrt(2) = (mq -pn)/nq) contradiction since sqrt(2) is itself irrational